OpenStudy (anonymous):

two point charges 8q and -2q are located at x=0 and x=L . the location of a point on the x axis at which the net electric field due to these two point charges is zero is (a) 4L (b)8L (c)L/4 (d)2L

OpenStudy (anonymous):

must be C

OpenStudy (anonymous):

reason please ?

OpenStudy (anonymous):

"on the x axis at which the net electric field due to these two point charges is zero "

OpenStudy (anonymous):

even the others are on the x axis

OpenStudy (anonymous):

Answer is a) 4L

OpenStudy (anonymous):

yep @srivatsa1995

OpenStudy (anonymous):

in which schul do u study @srivatsa1995

OpenStudy (anonymous):

kv iisc

OpenStudy (anonymous):

@shivam_bhalla how did u get 4L

OpenStudy (anonymous):

Reason : Net field at the point must be zero Assume r-->distance of the point where electric field is zero from origin. Therefore \[(k)(8q)/r^2 + k (-2q) / (r-2L)^2 = 0\] therefore \[(r/r-2L)^2=4\] Therefore \[(r/r-2L)=2\] \[r = 2r-4L\] Therefore r=4L

OpenStudy (anonymous):

why did u take r- 2l the whole square

OpenStudy (anonymous):

E-field of 8q =8q/(4 pi ε (x^2)) E-field of -2q =-2q/(4 pi ε ((L-x)^2)) total E-field =8q/(4 pi ε (x^2)) + -2q/(4 pi ε ((L-x)^2)) 8q/(4 pi ε (x^2)) + -2q/(4 pi ε ((L-x)^2))=0 8(L^2 -2xL +x^2)-2x^2 =0 3x^2 -8xL +4L^2=0 (3x-2L)(x-2L)=0 x=2L

OpenStudy (anonymous):